# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

### v1

@@ 4.2 Precise knowledge of standard deviation
We will thus consider the setting where the seller knows the precise standard deviation which means $\underaccent{\bar}{\sigma}=\bar{\sigma}$. For notational convenience, we will drop the bar notation and write $\sigma$ instead. We solve
$\sup_{p}\hat{R}(p),$ (17)
where it follows from Lemma 3.1 for $\underaccent{\bar}{\sigma}=\bar{\sigma}$ that
@@
with $\mathcal{P}(\mu,\sigma,\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and standard deviation $\sigma$. Figure LABEL:figmain1h displays the worst-case revenue for several risk levels. Observe that $\hat{R}(p)$ still consists of the two quadratic functions, pasted together, and that both functions present a candidate for the global maximum. However, compared with Figure LABEL:figmain1, there are some notable differences. First, the price range with non-degenerate prices increases with $\sigma$, now allowing robust prices that exceed the mean $\mu$. Second, the risk level from which the high price starts dominating the low price seems to be lower than in Figure LABEL:figmain1. And third, the plots in Figure LABEL:figmain1h show that for high prices, the revenue increases significantly with $\sigma$, which leads to premium pricing targeted at users with the highest valuations. We will substantiate these observations in a second pricing theorem, for which we introduce a third price function
$p^{*}_{h}(\sigma,\beta)$ $=\beta-\sqrt{\beta\left(\beta-\mu-\frac{\sigma^{2}}{\mu}\right)},$ (19)
which unlike $p^{*}_{l}(\sigma)$ and $p^{*}_{m}(\beta)$ depends on both $\sigma$ and $\beta$. This price function turns out pivotal in high-risk settings:
@@
Hence, $R_{3}(p)$ is concave and attains a unique maximum at $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Since $R_{3}(p)$ decreases in $[p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{2}]$, the maximum of $R_{3}(p)$ in $[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$ is attained in $\underaccent{\bar}{\upsilon}_{1}$ when $\underaccent{\bar}{\upsilon}_{1}=\max\{p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{1}\}$. ∎
As in the proof Proposition 4.1, we next want to rule out $\underaccent{\bar}{\upsilon}_{1}$ as candidate optimal price, but this now turns out more challenging. The next lemma states that indeed $\underaccent{\bar}{\upsilon}_{1}$ can never be the optimal price, and the proof using some advanced analysis is deferred to the appendix.
@@
Lemmas 4.1 and 4.2 together show that the optimal price is either $p^{*}_{l}(\sigma)$ or $p^{*}_{h}(\sigma,\beta)$. The remaining step is to show that the low (high) price is optimal for low (high) variance.
@@
since $p^{*}_{h}(\sigma,\beta)\geq 0$. Together these two statements prove that the threshold value $\sigma^{*}$ exists and is unique. Next, consider $\sigma\xrightarrow{}0$ and assume $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Then $p^{*}_{h}(\sigma,\beta)=\mu$, which is a contradiction, and $\hat{p}^{*}=p^{*}_{l}(\sigma)$. Furthermore, consider $\sigma\xrightarrow{}\sigma_{\textup{max}}=\sqrt{\mu(\beta-\mu)}$ and assume $p^{*}_{l}(\sigma)\in[0,\underaccent{\bar}{\upsilon}_{1}]$. Then $p^{*}_{l}(\sigma)=0$, which cannot be optimal, so $\hat{p}^{*}=p^{*}_{h}(\sigma,\beta)$. ∎
@@
5 Robust pricing with three prices
We proceed with the most general setting covered in this study, where instead of knowing an upper bound or exact value, the seller knows that the standard deviation is contained in an interval $[\underaccent{\bar}{\sigma},\bar{\sigma}]$. The width of this interval thus reflects how confident the seller is about the actual value $\sigma$, or equivalently, how accurate the seller can estimate $\sigma$. From Lemma 3.1 we see that the revenue maximization problem becomes $\sup_{p}\tilde{R}(p)$ with
Here is this paper’s most general and main result for robust monopoly pricing:
Theorem 5.1 (Three prices theorem).
Proof. When $\underaccent{\bar}{\sigma}=\bar{\sigma}$, we know from Proposition 4.2 that $p^{*}\in\{p^{*}_{l}(\bar{\sigma}),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\}$. For the remainder of the proof, assume $\underaccent{\bar}{\sigma}<\bar{\sigma}$. From Proposition 4.1 and 4.2, we know that $p^{*}_{l}(\bar{\sigma}),p^{*}_{m}(\beta),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ are the optimal prices of the unconstrained functions $R_{1}(p),R_{2}(p),R_{3}(p)$ respectively. What remains is to exclude the potential boundary optimal prices $\bar{\upsilon}_{1}$ and $\underaccent{\bar}{\upsilon}_{1}$.
@@ 5 Robust pricing with three prices
Next, we will show that $\underaccent{\bar}{\upsilon}_{1}$ cannot be the unique optimal price. Notice that for $\underaccent{\bar}{\upsilon}_{1}$ to be the unique global maximum, the ordering $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)<\underaccent{\bar}{\upsilon}_{1}<p^{*}_{m}(\beta)$ has to hold. However, $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\geq p^{*}_{m}(\beta)$, as both prices exhibit the form of $\beta-\sqrt{\beta(\beta-c)}$ with $c$ being $\mu$ for $p^{*}_{m}(\beta)$ and $\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu}$ for $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$. Hence, we can exclude $\underaccent{\bar}{\upsilon}_{1}$ as an optimal price candidate as well. ∎
Although Theorem 5.1 largely resolves the robust pricing problem, as it reduces the set of infinitely many prices to three prices only, it does not provide insight into how the optimal price responds to changes in market volatility $\sigma$ or market valuation bound $\beta$. We therefore build on Theorem 5.1, and the extreme cases covered in Theorems 4.1 and 4.2, to obtain a more detailed characterization of the optimal pricing strategy.
Theorem 5.1 tells us that the optimal price can be found by three pairwise comparisons, for which we shall introduce three functions. The function $f$ is the value of $\bar{\sigma}$ for which $R_{1}(p^{*}_{l}(\bar{\sigma}))=R_{2}(p^{*}_{m}(\beta))$ and corresponds with the threshold value $\bar{\sigma}$ that appeared in Proposition 4.1:
@@
We can now characterize the optimal price in terms of risk.
@@
The optimal robust price $p^{*}=\arg\sup_{p}\tilde{R}(p)$ is
where the risk threshold sets are given by
$\Sigma_{l}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq h(\bar{\sigma},\mu,\beta),\bar{\sigma}\leq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$ (31)
$\Sigma_{m}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq g(\mu,\beta),\bar{\sigma}\geq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$
Figure 4 visualizes Theorem 5.2. When either $\underaccent{\bar}{\sigma}$ or $\bar{\sigma}$ is held constant and does not coincide with the intersection of $\Sigma_{l}$ and $\Sigma_{m}$, varying the other parameter will result in at most two relevant prices, but never three. The high pricing strategy $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ becomes optimal when $\underaccent{\bar}{\sigma}$ and $\bar{\sigma}$ are both sufficiently large.
[…]
Figure 4: Three prices theorem for $\mu=0.5$ and $\beta=1$.
5.1 Comparison with optimal pricing
As is common in the pricing literature, we will now evaluate the performance of the robust pricing strategy with the optimal Bayesian strategy; see e.g. Azar and Micali (2012) and Chen et al. (2022a). The optimal Bayesian strategy is known to be fixed pricing, see Riley and Zeckhauser (1983), and achieves the optimal expected revenue
$\textup{OPT}(\mathbb{P})\mathrel{\mathop{:}}=\sup_{p}p\mathbb{P}(X\geq p).$ (32)
Markov’s inequality yields
$\mathbb{P}(X\geq p)\leq\frac{\mu}{p}\Leftrightarrow p\mathbb{P}(X\geq p)\leq\mu,$ (33)
and hence $\textup{OPT}(\mathbb{P})\leq\mu$. This is in fact a tight bound. To see this, let $p\in[\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu},\mu+{\bar{\sigma}^{2}}/{\mu}]$ and consider the two-point distribution $\mathbb{P}_{0}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ with atoms $\{0,p\}$ and associated probabilities $1-\mu/p$ and ${\mu}/{p}$, respectively, for which $p\mathbb{P}_{0}(X\geq p)=\mu$. Combining this bound with Theorem 5.2 yields, upon some rewriting, the following performance guarantee for robust pricing:
Corollary 5.1.
$\frac{p^{*}\mathbb{P}(X\geq p^{*})}{\textup{OPT}(\mathbb{P})}\geq\begin{cases}\rho_{l}(\mu,\bar{\sigma})\mathrel{\mathop{:}}=\frac{1}{2}\frac{\bar{\sigma}}{\mu}\kappa^{3}(\mu,\bar{\sigma})\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{l},\\ \rho_{m}(\mu,\beta)\mathrel{\mathop{:}}=\frac{2}{\mu}(\beta-\sqrt{\beta(\beta-\mu)})-1\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{m},\\ \rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)\mathrel{\mathop{:}}=2(1-\sqrt{1-\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}})-\tfrac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{h}.\end{cases}$ (34)
[…]
Figure 5: Corollary 5.1 with $\mu=0.5$ and $\beta=1$.
Figure 5 illustrates Corollary 5.1. When the seller has precise information about the standard deviation, there occurs a U-shaped performance guarantee, aligned with the U-shaped revenue curves found in monopolistic markets Johnson and Myatt (2006), Sun (2012) discussed in the introduction. In our case, this U-shape occurs as the seller switches from low to high pricing, changing the curve from being decreasing to increasing. More generally, we have found that the U-shape occurs when the seller has precise enough (not perfect per se) information on the standard deviation. When the seller only knows an upper bound, the U-shape is replaced with a function that first decreases and then remains constant, see Figure 5. We substantiate these numerical insights with the following result:
Corollary 5.2 (Properties of price and revenue).
The optimal price and performance guarantee thus both move in the same direction in response to changing mean, support and variance. In volatile markets, the essence of successful pricing lies in its ability to adjust prices in accordance with uncertainty levels, ensuring higher prices with increasing volatility. Corollary 5.2 shows that out of the three potential optimal price functions, only the high price is positively correlated with variance, resulting in a performance guarantee that also increases with variance.
5.2 Effect of bundling
The existing literature considering mean-variance information described a pricing policy with only one robust price Azar and Micali (2012), Chen et al. (2022a), the low price in our pricing policy. We also include support information, and find two and at most three price functions. Our new pricing policy that adjusts the price to market volatility is transparent and simple with as main strength a high price responding to high uncertainty, leading to high revenue. If the goal is to use only one, instead of two or three prices, the seller could opt for bundling. The seller then sets a price for a bundle of goods, so that customers buy either the bundle or nothing. A big theme in the pricing literature centers around the question whether bundling is better than separate sales, and in general there is no clear or straightforward answer to that question Bakos and Brynjolfsson (1999), Hart and Nisan (2017). We will use the result for the single product to assess this bundle product pricing. Assuming valuations are i.i.d., the revenue of selling a bundle of $m$ goods with valuations $X_{1},\ldots,X_{m}$ for a bundle price $p_{(m)}$ has expected revenue
$p_{(m)}\cdot\mathbb{P}\Big{(}\sum_{i=1}^{m}X_{i}\geq p_{(m)}\Big{)}.$ (35)
Turning to the limited information setting, the single-product information about the mean, variance and maximal valuation translates into information for the bundle by multiplying all three characteristics with the number of goods $m$. Notice that a bundle of $m$ i.i.d. goods, each with mean $\mu$, standard deviation $\sigma\in[\underaccent{\bar}{\sigma},\bar{\sigma}]$ and maximal valuation $\beta$ can be viewed as selling a single good with as valuation the average of all $m$ valuations, so with mean $\mu$, maximal value $\beta$ and $\sigma\in[\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\frac{\bar{\sigma}}{\sqrt{m}}]$. As a result, for any $p<\mu$ and large enough $m$, the tail bound (6) becomes
$\inf_{\mathbb{P}\in\mathcal{P}\left(\mu,\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\frac{\bar{\sigma}}{\sqrt{m}},\beta\right)}\mathbb{P}(X\geq p)=\frac{(\mu-p)^{2}}{(\mu-p)^{2}+\tfrac{\bar{\sigma}^{2}}{m}}$ (36)
and the expected worst-case revenue is maximized by the low price $p^{*}_{l}(\mu,\frac{\sigma}{\sqrt{m}},\beta)$, yielding an optimal price and revenue that approach $\mu$ when $m$ tends to infinity. This result also follows from Chen et al. (2022a), Giannakopoulos et al. (2023) who consider robust pricing when the mean and variance (and not the bound on support) are known.
For ease of exposition, consider $\sigma=\underaccent{\bar}{\sigma}=\bar{\sigma}$, in which case the low price is guaranteed to be optimal when $\bar{\upsilon}_{1}\geq p^{*}_{h}(\mu,\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\beta)$, so for $m\geq m^{*}$ with $m^{*}$ the solution to
$\mu-\frac{\sigma^{2}}{m^{*}(\beta-\mu)}=\beta-\sqrt{\beta\left(\beta-\mu-\frac{\sigma^{2}}{m^{*}\mu}\right)}$ (37)
given by
$m^{*}=\left(\frac{\left(1+\sqrt{\frac{\beta+3\mu}{\beta-\mu}}\right)\beta+2\mu}{2\mu^{2}(\beta-\mu)}\right)\sigma^{2}.$ (38)
While this result is not tight, it shows that the bundle size needed to guarantee that the low price $p^{*}_{l}(\mu,\frac{\sigma}{\sqrt{m}})$ is optimal grows proportionally with the valuation variance, and that bundling in this setting with i.i.d. valuations thus leads to low pricing when the bundle is large enough.

### v2

@@ 4.2 Precise knowledge of standard deviation
Next, consider the setting where the seller knows the precise standard deviation which means $\underaccent{\bar}{\sigma}=\bar{\sigma}$. For notational convenience, we will drop the bar notation and write $\sigma$ instead. We solve $\sup_{p}\hat{R}(p)$, where it follows from Lemma 3.1 for $\underaccent{\bar}{\sigma}=\bar{\sigma}$ that
@@
with $\underaccent{\bar}{\upsilon}_{1}$ and $\underaccent{\bar}{\upsilon}_{2}$ as in (5) and $\mathcal{P}(\mu,\sigma,\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and standard deviation $\sigma$. Figure LABEL:figmain1h displays the worst-case revenue for several risk levels. Observe that $\hat{R}(p)$ still consists of the two quadratic functions, pasted together, and that both functions present a candidate for the global maximum. However, compared with Figure LABEL:figmain1, there are some notable differences. First, the price range with non-degenerate prices increases with $\sigma$, now allowing robust prices that exceed the mean $\mu$. Second, the risk level from which the high price starts dominating the low price seems to be lower than in Figure LABEL:figmain1. And third, the plots in Figure LABEL:figmain1h show that for high prices, the revenue increases significantly with $\sigma$, which leads to premium pricing targeted at users with the highest valuations.
Taken together, this gives the following pricing result:
@@
Hence, $R_{3}(p)$ is concave and attains a unique maximum at $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Since $R_{3}(p)$ decreases in $p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{2}]$, the maximum of $R_{3}(p)$ in $\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$ is attained in $\underaccent{\bar}{\upsilon}_{1}$ when $\underaccent{\bar}{\upsilon}_{1}=\max\{p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{1}\}$. ∎
As in the proof of Proposition 4.1, we want to rule out $\underaccent{\bar}{\upsilon}_{1}$ as candidate optimal price, but this now turns out more challenging. The next lemma states that indeed $\underaccent{\bar}{\upsilon}_{1}$ can never be the optimal price, and the proof using some advanced analysis is deferred to the appendix.
@@
Lemmas 4.1 and 4.2 together show that the optimal price is either $p^{*}_{l}(\sigma)$ or $p^{*}_{h}(\sigma,\beta)$. The remaining step is to show that the low (high) price is optimal for low (high) standard deviation.
@@
since $p^{*}_{h}(\sigma,\beta)\geq 0$. Together these two statements prove that the threshold value $\sigma^{*}$ exists and is unique.
Next, consider $\sigma\xrightarrow{}0$ and assume $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Then $p^{*}_{h}(\sigma,\beta)=\mu$, which is a contradiction, and $\hat{p}^{*}=p^{*}_{l}(\sigma)$. Furthermore, consider $\sigma\xrightarrow{}\sigma_{\textup{max}}=\sqrt{\mu(\beta-\mu)}$ and assume $p^{*}_{l}(\sigma)\in[0,\underaccent{\bar}{\upsilon}_{1}]$. Then $p^{*}_{l}(\sigma)=0$, which cannot be optimal, so $\hat{p}^{*}=p^{*}_{h}(\sigma,\beta)$. ∎
@@
4.3 Proof of Theorem 3.1
When $\underaccent{\bar}{\sigma}=\bar{\sigma}$, we know from Proposition 4.2 that $p^{*}\in\{p^{*}_{l}(\bar{\sigma}),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\}$. For the remainder of the proof, assume $\underaccent{\bar}{\sigma}<\bar{\sigma}$. From Proposition 4.1 and 4.2, we know that $p^{*}_{l}(\bar{\sigma}),p^{*}_{m}(\beta),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ are the optimal prices of the unconstrained functions $R_{1}(p),R_{2}(p),R_{3}(p)$ respectively. What remains is to exclude the potential boundary optimal prices $\bar{\upsilon}_{1}$ and $\underaccent{\bar}{\upsilon}_{1}$.
@@
Next, we will show that $\underaccent{\bar}{\upsilon}_{1}$ cannot be the unique optimal price. Notice that for $\underaccent{\bar}{\upsilon}_{1}$ to be the unique global maximum, the ordering $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)<\underaccent{\bar}{\upsilon}_{1}<p^{*}_{m}(\beta)$ has to hold. However, $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\geq p^{*}_{m}(\beta)$, as both prices exhibit the form of $\beta-\sqrt{\beta(\beta-c)}$ with $c$ being $\mu$ for $p^{*}_{m}(\beta)$ and $\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu}$ for $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$. Hence, we can exclude $\underaccent{\bar}{\upsilon}_{1}$ as an optimal price candidate as well.
We next build on the insight that there are three prices, and the extreme cases covered in Theorems 4.1 and 4.2, to obtain a more detailed characterization of the optimal pricing strategy based on three pairwise comparisons, for which we shall introduce three functions. The function $f$ is the value of $\bar{\sigma}$ for which $R_{1}(p^{*}_{l}(\bar{\sigma}))=R_{2}(p^{*}_{m}(\beta))$ and corresponds with the threshold value $\bar{\sigma}$ that appeared in Proposition 4.1:
@@
We can now characterize the optimal price in terms of risk. The optimal robust price $p^{*}=\arg\sup_{p}\tilde{R}(p)$ is then as in (15) with the risk threshold sets given by
$\Sigma_{l}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq h(\bar{\sigma},\mu,\beta),\bar{\sigma}\leq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$
$\Sigma_{m}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq g(\mu,\beta),\bar{\sigma}\geq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$ (29)
4.4 Proof of Theorem 3.2
We now show that as the bounds of the standard deviation interval increase, a discontinuous jump eventually occurs in which we transition from the low price $p^{*}_{l}(\bar{\sigma})$ to a higher price, either $p^{*}_{m}(\beta)$ or $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$.
We begin by reformulating Theorem 3.2, originally stated in fairly loose terms, into a more quantifiable and precise form. We allow $\underaccent{\bar}{\sigma}$ and $\bar{\sigma}$ to increase in a general manner by letting $\bar{\sigma}$ increase while imposing $\underaccent{\bar}{\sigma}=\Phi(\bar{\sigma})$, where $\Phi$ is any non-decreasing function of $\bar{\sigma}$ constrained by $\Phi(\bar{\sigma})\leq\bar{\sigma}$. Special cases include the standard deviation as an upper bound when $\Phi(\bar{\sigma})=0$ and as exact when $\Phi(\bar{\sigma})=\bar{\sigma}$.
@@
When $\bar{\sigma}<f(\mu,\beta)$ and $\Phi(\bar{\sigma})<h(\bar{\sigma},\mu,\beta)$, then
$\tilde{R}(p^{*}_{l}(\bar{\sigma}))<\max\{\tilde{R}(p^{*}_{m}(\beta)),\tilde{R}(p^{*}_{h}(\Phi(\bar{\sigma}),\beta))\},$
and when $\bar{\sigma}>f(\mu,\beta)$ or $\Phi(\bar{\sigma})>h(\bar{\sigma},\mu,\beta)$, then $\tilde{R}(p^{*}_{l}(\bar{\sigma}))>\max\{\tilde{R}(p^{*}_{m}(\beta)),\tilde{R}(p^{*}_{h}(\Phi(\bar{\sigma}),\beta))\}$.
Proof. First, since $R_{1}(p)$ is maximal and continuously differentiable in $p^{*}_{l}(\bar{\sigma})$, we can employ the envelope theorem to show that
since $p^{*}_{l}(\bar{\sigma})\geq 0$. Second,
$\frac{\partial R_{2}(p^{*}_{m}(\beta))}{\partial\bar{\sigma}^{2}}=0,$ (31)
as $R_{2}(p^{*}_{m}(\beta))$ does not depend on $\bar{\sigma}$. Finally, since $R_{3}(p)$ is maximal and continuously differentiable in $p^{*}_{h}(\Phi(\bar{\sigma}),\beta)$, the envelope theorem yields
$\frac{\partial R_{3}(p^{*}_{h}(\Phi(\bar{\sigma}),\beta))}{\partial\Phi(\bar{\sigma})^{2}}=\frac{\partial R_{3}(p)}{\partial\Phi(\bar{\sigma})^{2}}\biggr{\rvert}_{p=p^{*}_{h}(\Phi(\bar{\sigma}),\beta)}=\frac{p^{*}_{h}(\Phi(\bar{\sigma}),\beta)}{\beta(\beta-p^{*}_{h}(\Phi(\bar{\sigma}),\beta))}\geq 0,$ (32)
since $p^{*}_{h}(\Phi(\bar{\sigma}),\beta)\geq 0$. Hence, since $\Phi(\bar{\sigma})$ is increasing in $\bar{\sigma}$, we also have that $R_{3}(p^{*}_{h}(\Phi(\bar{\sigma}),\beta))$ is increasing in $\bar{\sigma}$.
Next, consider $(\Phi(\bar{\sigma}),\bar{\sigma})=(0,0)$. Then $\bar{\upsilon}_{1}=\underaccent{\bar}{\upsilon}_{1}=\underaccent{\bar}{\upsilon}_{2}=\mu$, thus $p^{*}=p^{*}_{l}(\bar{\sigma})$. Furthermore, consider $(\Phi(\bar{\sigma}),\bar{\sigma})=(\Phi(\sigma_{\textup{max}}),\sigma_{\textup{max}})=(\Phi(\sqrt{\mu(\beta-\mu)}),\sqrt{\mu(\beta-\mu)})$. Then $\underaccent{\bar}{\upsilon}_{1}=0$, thus $p^{*}\in\{p^{*}_{m}(\beta),p^{*}_{h}(\Phi(\bar{\sigma}),\beta)\}$. ∎
Theorem 4.1 displays the robustness of our main insight: even when the standard deviation lies within an interval $\underaccent{\bar}{\sigma},\bar{\sigma}]$, an abrupt shift from a lower to a higher price still occurs. However, in contrast to the specific cases where the standard deviation is an upper bound or exact, the general scenario may exhibit two discontinuous jumps depending on the choice of $\Phi$. Notably, these shifts always follow a clear progression from the low price to the middle price, and then to the high price. This key insight follows from the proof of Theorem 4.1, which reveals that $R_{2}(p^{*}_{m}(\beta))$ is independent of $\bar{\sigma}$, while $R_{3}(p^{*}_{h}(\Phi(\bar{\sigma}),\beta))$ is increasing in $\bar{\sigma}$.